OpenStudy (anonymous):

in the xy-plane, the line x + y=k, where k is a constant, is tangent to the graph of y=x^2+3x+1. What is the value of k?

OpenStudy (dumbcow):

slope of tangent line = dy/dx = 2x+3 k is y_intercept of tangent line 2x+3 = -1 x = -2 y = (-2)^2+3(-2)+1 = -1 tangent line passes through point (-2,-1) k = x+y k = -2 -1 k = -3

OpenStudy (anonymous):

thank you! why did you set 2x+3= -1 though? where did the -1 come from

OpenStudy (dumbcow):

oh sorry its the given slope of tangent line x+y=k --> y = -x+k

OpenStudy (anonymous):

OH! of course! haha thanks so much! i get it now

OpenStudy (dumbcow):

welcome